# Structured Reinforcement Learning for Combinatorial Decision-Making

### v1

@@ Abstract
Reinforcement learning (RL) is increasingly applied to real-world problems involving complex and structured decisions, such as routing, scheduling, and assortment planning. These settings challenge standard RL algorithms, which struggle to scale, generalize, and exploit structure in the presence of combinatorial action spaces. We propose Structured Reinforcement Learning (SRL), a novel actor-critic framework that embeds combinatorial optimization layers into the actor neural network. We enable end-to-end learning of the actor via Fenchel-Young losses and provide a geometric interpretation of SRL as a primal-dual algorithm in the dual of the moment polytope. Across six environments with exogenous and endogenous uncertainty, SRL matches or surpasses the performance of unstructured RL and imitation learning on static tasks and improves over these baselines by up to 92% on dynamic problems, with improved stability and convergence speed. [^1] Our code is available at https://github.com/tumBAIS/Structured-RL.
@@ 1 Introduction
[see Appendix A for details] We consider MDPs with states $s\in\mathcal{S}$, actions $a\in\mathcal{A}(s)\subset\mathbb{R}^{d(s)}$, rewards $r$, and transitions $\mathbb{P}(s^{\prime},r\mid s,a)$. In combinatorial MDPs, $\mathcal{A}(s)$ is given by the feasible set of a combinatorial problem; we denote its convex hull by $\mathcal{C}(s):=\operatorname{conv}(\mathcal{A}(s))$, also called moment polytope. The goal is to find a policy $\bar{\pi}$ maximizing the expected discounted return:
$\bar{\pi}\in\arg\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right].$
The limitations of standard RL algorithms in solving such Combinatorial MDPs (C-MDPs) render them insufficient to solve many industrial problems. To overcome these drawbacks, we propose Structured Reinforcement Learning (SRL)—a novel actor-critic RL paradigm that embeds combinatorial optimization (CO)-layers into neural actors, enabling to exploit the underlying problem structure.
@@
Prior approaches to solving C-MDPs include handcrafted decision rules (e.g., Liyanage and Shanthikumar, 2005; Huber et al., 2019) and predict-and-optimize algorithms (e.g., Alonso-Mora et al., 2017; Bertsimas and Kallus, 2020). While the former fail to model complex dynamics or constraints, the latter typically rely on imitation learning and separate prediction from optimization, which impairs performance in dynamic settings (cf., Enders et al., 2023). In contrast, CO-augmented Machine Learning (COaML)-pipelines integrate combinatorial optimization directly into model architectures, allowing end-to-end learning (e.g., Parmentier, 2022; Dalle et al., 2022). Here, the key challenge is to differentiate through the CO-layer. Existing approaches include problem-specific relaxation schemes (e.g., Vlastelica et al., 2020), and more general strategies, e.g., using Fenchel-Young losses (e.g., Blondel et al., 2020; Berthet et al., 2020). The latter allow continuous training over discrete structures and are increasingly used for differentiating COaML-pipelines (e.g., Dalle et al., 2022). In C-MDPs, prior work either uses offline expert imitation (e.g., Baty et al., 2024) or treats the CO-layer as an action mask in unstructured RL (e.g., Hoppe et al., 2024). The former requires access to expert solutions, while the latter often leads to unstable gradients. Imitation approaches also suffer from insufficient exploration in multi-stage problems.
SRL builds on classical policy-gradient algorithms, e.g., REINFORCE, PPO, and SAC (cf. Williams, 1992; Haarnoja et al., 2018), which perform well in conventional MDPs but are challenged by the size and structure of C-MDP action spaces (Hildebrandt et al., 2023). Problem-specific neural network designs often lack generalizability across applications and struggle in dynamic contexts (e.g., Bello et al., 2017; Dai et al., 2017), while neural improvement methods rely on existing initial solutions (e.g., Yuan et al., 2022; Hottung and Tierney, 2022). Value-based methods often assume decomposable critics (e.g., Xu et al., 2018; Liang et al., 2022), limiting applicability in structured domains. SRL is also related to offline RL, which frequently relies on imitation learning (e.g., Figueiredo Prudencio et al., 2024), but differs by operating online and updating from on-policy targets.
@@ 2.1 Structured Reinforcement Learning
Consider a C-MDP as detailed in Appendix A, where $\mathcal{A}(s)$ denotes the feasible action set in state $s$. Deep RL algorithms typically use neural networks to represent the policy as a distribution over the action space. In combinatorial settings, a neural network may struggle to encode a distribution on the exponentially-large action space $\mathcal{A}(s)$. To overcome this challenge, we adopt a CO-augmented Machine Learning-pipeline as the actor architecture. As illustrated in Figure 1, this architecture combines a statistical model $\varphi_{w}$ with a combinatorial optimizer $f$. The statistical model $\varphi_{w}$, usually a neural network parameterized by weights $w$, observes contextual information provided by the C-MDP state $s$. Using this information, the model estimates a latent score vector $\theta$ with dimension $d$. The combinatorial optimizer $f$ uses $\theta$ as coefficients of a linear objective function and generates an action $a$ by solving the CO problem
$a=f(\theta,s)\in\operatorname*{argmax}_{\tilde{a}\in\mathcal{A}(s)}\>\langle\theta|\tilde{a}\rangle$
@@
In the COaML-pipeline, the statistical model $\varphi_{w}$ encodes contextual information, which enables generalizing across states, capturing variable dependencies, and anticipating C-MDP dynamics. The CO-layer $f$ enforces combinatorial feasibility, facilitates a structured exploration of the action space, and improves scalability by mapping score vectors $\theta$ to potentially high-dimensional action spaces $\mathcal{A}(s)$. For a detailed introduction to COaML-pipelines and a motivating example, we refer to Appendix B.
@@
        Update actor using $\mathcal{L}_{\Omega}(\theta;\widehat{a})$triangleright using a second perturbation
@@
Originally proposed for imitation learning (Blondel et al., 2020; Berthet et al., 2020), the Fenchel-Young loss has become an established loss function for the end-to-end training of COaML-pipelines (e.g., Parmentier and T’Kindt, 2023; Baty et al., 2024). We adopt it in SRL since it is convex in the output of the statistical model $\theta=\varphi_{w}(s)$ and differentiable with respect to the latter, while leveraging the structure of the CO-layer $f$. We refer to Appendix C for a full description of the Fenchel-Young loss, and briefly recall its definition here. Given a regularization function $\Omega:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ and its Fenchel conjugate $\Omega^{*}$, the Fenchel-Young loss $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ generated by $\Omega$ is defined over $\operatorname{dom}(\Omega^{*})\times\operatorname{dom}(\Omega)$ as
$\mathcal{L}_{\Omega}(\theta;\widehat{a}):=\Omega^{*}(\theta)+\Omega(\widehat{a})-\langle\theta|\widehat{a}\rangle=\sup_{a\in\operatorname{dom}(\Omega)}\big{(}\langle\theta|a\rangle-\Omega(a)\big{)}-\big{(}\langle\theta|\widehat{a}\rangle-\Omega(\widehat{a})\big{)}.$ (1)
@@
with temperature parameter $\tau$, controlling the sharpness of the softmax. If we dcrease $\tau$, $\widehat{a}$ approaches the greedy action $\operatorname*{argmax}_{a^{\prime}}\left(Q_{\psi_{\beta}}(s,a^{\prime})\right)$. If we increase $\tau$, $\widehat{a}$ approaches the uniform average $\frac{1}{m}\sum_{a^{\prime}}a^{\prime}$. Note that $\widehat{a}\in\mathcal{A}(s)$ does not have to hold, as it is not executed in the environment, but only serves as a training signal for the Fenchel-Young loss.
Figure 3: Schematic representation of SRL update step: unperturbed action $a$ (left), perturbed actions and target action $\widehat{a}$ (middle), Fenchel-Young loss $\mathcal{L}_{\Omega}(\theta;\widehat{a})$ (right).
@@
To ensure sufficient exploration of the state space $\mathcal{S}$, we introduce stochasticity into the forward pass during training: we perturb the score vector $\theta$ using Gaussian noise $Z\sim\mathcal{N}(\theta,\sigma_{f})$ with exploration standard deviation $\sigma_{f}$, sample one perturbed vector $\eta$, and select an action $a=f(\eta,s)$. We promote exploration of the combinatorial action space $\mathcal{A}(s)$ by using a large enough perturbation standard deviation $\sigma_{b}$ and a sufficient number of samples $m$. These measures encourage diverse environment interactions, essential for robust policy learning.
@@ 2.2 Geometrical insights
Policies as solutions of regularized optimization over the distribution simplex.
@@
A sampling-based primal-dual algorithm for the actor update.
@@
Here, $\Omega^{*}$ is the Fenchel-conjugate of $\Omega$, $\nabla\Omega^{*}$ the gradient of $\Omega^{*}$, and $\partial\Omega$ the sub-differential of $\Omega$. The following proposition, proved in Appendix D, shows that the static version of Algorithm 1 can be seen as a variant of the primal-dual algorithm presented in Bouvier et al. (2025), enhancing it with an additional sampling step.
@@
where $\delta_{a}$ is the Dirac distribution on $a$, $\Omega_{\Delta}$ is the negentropy, and $\Omega_{\varepsilon,\Delta}$ is the conjugate of the sparse perturbation, both detailed in Appendix D. Since $\hat{q}_{m}^{(t+\frac{1}{2})}$ is sparse by design, we discuss the definition of gradients and sub-gradients at the (relative) boundary of the domains in Appendix D.
In Equations (6), for convenience in the implementation, we involve two distinct regularization functions on the distribution simplex $\Delta^{\mathcal{A}}$, detailed in Appendix D. The main difference between Equations (5) and Equations (6) is the sampling step for the primal update. Recall that involving the critic in Equation (5a) may be intractable. Indeed, the critic function may be highly nonlinear, and the resulting nonlinear combinatorial optimization problem may be intractable. Instead, Equation (6) is based on a sampling step, which only requires $m$ evaluations of the critic, which is tractable even when optimizing it with respect to $a\in\mathcal{A}(s)$ is not. In both primal-dual algorithms, the dual update (of $\theta$) is equivalent to solving a convex optimization problem, precisely minimizing a Fenchel-Young loss generated by $\Omega_{\mathcal{C}}$. It is very convenient in practice, since the weights $w$ of our actor policy can be updated via stochastic gradient descent, although it relies on a piecewise constant CO-layer $f$.
@@ 3 Numerical studies
We evaluate SRL across six environments that reflect industrial applications with large combinatorial action spaces. Appendix E describes the experimental setup, and Appendix F details the environments. We first consider three static environments—common industrial benchmarks from Dalle et al. (2022)—namely, a Warcraft Shortest Paths Problem, a Single Machine Scheduling Problem, and a Stochastic Vehicle Scheduling Problem. As discussed in Appendix G, SRL matches the performance of SIL, while relying solely on access to a (black-box) cost function and without requiring the expert knowledge necessary for SIL. These results underline the versatility and broader potential of SRL, even though we designed it primarily for dynamic settings. We now discuss our findings for the dynamic environments in more detail.
The dynamic environments model online decision-making in C-MDPs. We consider: i) A Dynamic Vehicle Scheduling Problem (DVSP), based on the Dynamic Vehicle Routing Problem introduced by Kool et al. (2022); Baty et al. (2024). This problem with exogenous uncertainty requires computing cost-minimizing routes to service spatio-temporally distributed demands revealed over time. ii) A Dynamic Assortment Problem (DAP), adapted from Dulac-Arnold et al. (2016) andChen et al. (2020). This problem with endogenous uncertainty involves selecting revenue-maximizing item assortments, where customer choices follow a multinomial logit model and item features evolve based on past decisions. iii) A Gridworld Shortest Paths Problem (GSPP), inspired by gridworld and robotic control tasks (Chandak et al., 2019; Zhang et al., 2020). This problem with endogenous uncertainty requires the agent to find cost-minimizing paths to a moving target, with costs influenced by prior actions.
@@
We compare SRL against two baselines: SIL, a structured imitation learning approach, and Proximal Policy Optimization (PPO), an unstructured RL algorithm. All algorithms use identical COaML-pipelines, and SRL and PPO also share the same critic architectures to ensure a fair comparison. We select SIL due to its methodological proximity to SRL and its strong performance in combinatorial settings (Baty et al., 2024; Jungel et al., 2024). We include PPO for its stability and compatibility with our pipeline architecture (Schulman et al., 2017), noting that alternatives such as Soft Actor-Critic would require substantial modifications to the COaML-pipelines and critic design. As performance references, we include two additional baselines: an expert policy and a greedy policy. In the static DVSP, the expert has access to the complete problem instance and thus represents the offline optimum. In contrast, in the dynamic DAP and GSPP, traceability constraints limit information access, and the expert corresponds to the best possible online policy, given the sequential nature of the decision process. We train all algorithms using the same number of episodes, employing environment-specific train/validation/test splits. We tune hyperparameters per algorithm and environment, using the PPO-optimized episode numbers consistently across methods. Each algorithm is retrained using ten random seeds. Appendix E provides further details on the experimental setup and baselines.
@@
We present results for the dynamic environments in Figure 4 and Figure 5. While SRL performs comparably to SIL in the DVSP, it outperforms SIL in the DAP by 8% and in the GSPP by 78%, even surpassing the online optimum by 79% in the latter. In the DAP, the online optimum remains above the performance of SRL, but SIL fails to reach it. These results highlight two key limitations of imitation learning: i) its performance is bounded by that of the expert policy; and ii) it lacks exploration to learn policies that enable escape from suboptimal states. In contrast, PPO consistently underperforms across all environments, struggling to reach greedy policies—SRL outperforms it by 16% in the DVSP, 77% in the DAP, and 92% in the GSPP. This poor performance highlights the challenges faced by unstructured RL in combinatorial action spaces. Overall, these performance gains show the superiority of SRL over SIL and PPO.
@@
Stability metrics in Table 1 further explain these trends. PPO shows the highest variance across all environments—up to 80$\times$ higher than SRL and 40$\times$ higher than SIL—reflecting known limitations of unstructured RL in combinatorial settings (e.g., Enders et al., 2023; Hoppe et al., 2024). In contrast, SRL and SIL exhibit consistently low variance, underscoring the robustness of structured approaches.

### v2

@@ Abstract
Reinforcement learning (RL) is increasingly applied to real-world problems involving complex and structured decisions, such as routing, scheduling, and assortment planning. These settings challenge standard RL algorithms, which struggle to scale, generalize, and exploit structure in the presence of combinatorial action spaces. We propose Structured Reinforcement Learning (SRL), a novel actor-critic paradigm that embeds combinatorial optimization-layers into the actor neural network. We enable end-to-end learning of the actor via Fenchel-Young losses and provide a geometric interpretation of SRL as a primal-dual algorithm in the dual of the moment polytope. Across six environments with exogenous and endogenous uncertainty, SRL matches or surpasses the performance of unstructured RL and imitation learning on static tasks and improves over these baselines by up to 92% on dynamic problems, with improved stability and convergence speed. [^1] Our code is available at https://github.com/tumBAIS/Structured-RL.
@@ 1 Introduction
We consider MDPs with states $s\in\mathcal{S}$, actions $a\in\mathcal{A}(s)\subset\mathbb{R}^{d(s)}$, rewards $r$, and transition probabilities $\mathbb{P}(s^{\prime},r\mid s,a)$ with next state $s^{\prime}$. In combinatorial MDPs, $\mathcal{A}(s)$ is the set of feasible solutions of a combinatorial problem. We denote its convex hull as the moment polytope $\mathcal{C}(s):=\operatorname{conv}(\mathcal{A}(s))$. As in stochastic optimization (Bertsekas and Shreve, 1996), we use a latent noise variable $\xi\in\Xi$ with probability $p(\xi\mid s,a)$, such that the transition to $(s^{\prime},r)$ given $(s,a,\xi)$ is deterministic. We distinguish exogenous noise, where the distribution of $\xi$ does not depend on $a$, and endogenous noises, where the distribution of $\xi$ depends on $a$.
$\bar{\pi}\in\arg\max_{\pi}\mathbb{E}_{\pi,\mathbb{P}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$
$Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\mathbb{P}}\left[r_{t}\right]+\gamma\>\mathbb{E}_{\mathbb{P}}\left[\max_{\tilde{a}_{t+1}}Q^{\pi}(s_{t+1},\tilde{a}_{t+1})\right],$
using expectations with respect to $\mathbb{P}(s^{\prime},r\mid s,a)$.
The limitations of standard RL algorithms in solving such Combinatorial MDPs (C-MDPs) render them insufficient to solve many industrial problems. To overcome these drawbacks, we propose Structured Reinforcement Learning (SRL) – a novel actor-critic RL paradigm that embeds combinatorial optimization (CO)-layers into neural actors, enabling to exploit the underlying problem structure.
@@
Prior approaches to solving C-MDPs include handcrafted decision rules (e.g., Liyanage and Shanthikumar, 2005; Huber et al., 2019) and predict-and-optimize algorithms (e.g., Alonso-Mora et al., 2017; Bertsimas and Kallus, 2020). While the former fail to model complex dynamics or constraints, the latter typically rely on imitation learning and separate prediction from optimization, which impairs performance in dynamic settings (cf., Enders et al., 2023). In contrast, CO-augmented Machine Learning (COAML)-pipelines integrate combinatorial optimization directly into model architectures, allowing end-to-end learning (e.g., Parmentier, 2022; Dalle et al., 2022). Here, the key challenge is to differentiate through the CO-layer. Existing approaches include problem-specific relaxation schemes (e.g., Vlastelica et al., 2020), and more general strategies, e.g., using Fenchel-Young losses (e.g., Blondel et al., 2020; Berthet et al., 2020). The latter allow continuous training over discrete structures and are increasingly used for differentiating COAML-pipelines (e.g., Dalle et al., 2022). In C-MDPs, prior work either uses offline expert imitation (e.g., Baty et al., 2024; Jungel et al., 2024) or treats the CO-layer as an action mask in unstructured RL (e.g., Hoppe et al., 2024; Woywood et al., 2025). The former requires access to expert solutions, while the latter often leads to unstable gradients. Imitation approaches also suffer from insufficient exploration in multi-stage problems.
SRL builds on classical policy-gradient algorithms, e.g., REINFORCE, PPO, and SAC (cf., Williams, 1992; Haarnoja et al., 2018), which perform well in conventional MDPs but are challenged by the size and structure of C-MDP action spaces (Hildebrandt et al., 2023). Problem-specific neural network designs often lack generalizability across applications and struggle in dynamic contexts (e.g., Bello et al., 2017; Dai et al., 2017), while neural improvement methods rely on existing initial solutions (e.g., Yuan et al., 2022; Hottung and Tierney, 2022). Value-based methods often assume decomposable critics (e.g., Xu et al., 2018; Liang et al., 2022), limiting applicability in structured domains. Alternative approaches transform a task-specific Q-network into a mixed-integer linear program, which potentially increases solution time and limits the range of usable network architectures (e.g., Xu et al., 2025). SRL is also related to offline RL, which frequently relies on imitation learning (e.g., Figueiredo Prudencio et al., 2024), but differs by operating online and updating from on-policy targets.
@@ 2.1 Structured Reinforcement Learning
Consider a C-MDP as detailed in Problem 1, where $\mathcal{A}(s)$ denotes the feasible action set in state $s$. Deep RL algorithms typically use neural networks to represent the policy as a distribution over the action space. In combinatorial settings, a neural network may struggle to encode a distribution on the exponentially-large action space $\mathcal{A}(s)$. To overcome this challenge, we adopt a CO-augmented Machine Learning-pipeline as the actor architecture. As illustrated in Figure 1, this architecture combines a statistical model $\varphi_{w}$ with a combinatorial optimizer $f$. The statistical model $\varphi_{w}$, usually a neural network parameterized by weights $w$, observes contextual information provided by the C-MDP state $s$. Using this information, the model estimates a latent score vector $\theta$ with dimension $d$. The combinatorial optimizer $f$ uses $\theta$ as coefficients of a linear objective function and generates an action $a$ by solving the CO problem
$f(\theta,s):a=\operatorname*{argmax}_{\tilde{a}\in\mathcal{A}(s)}\>\langle\theta|\tilde{a}\rangle$
@@
In the COAML-pipeline, the statistical model $\varphi_{w}$ encodes contextual information, which enables generalizing across states, capturing variable dependencies, and anticipating C-MDP dynamics. The CO-layer $f$ enforces combinatorial feasibility, facilitates a structured exploration of the action space, and improves scalability by mapping score vectors $\theta$ to potentially high-dimensional action spaces $\mathcal{A}(s)$. COAML-pipelines have been used successfully to address combinatorial problems using imitation learning, highlighting their suitability for C-MDPs (e.g., Parmentier, 2022; Baty et al., 2024). For a detailed introduction to COAML-pipelines and a motivating example, we refer to Appendix A.
@@
   Update actor $\mathcal{L}_{\Omega}(\theta;\widehat{a}) \triangleright using$ a second perturbation
@@
[…]
$\mathcal{L}_{\Omega}(\theta;\widehat{a})=\max_{a\in\mathcal{A}(s)}\theta^{\top}a-\theta^{\top}\widehat{a}$. (1)
[…]
$\mathcal{L}_{\Omega}(\theta;\widehat{a})=\mathbb{E}\left[\max_{a\in\mathcal{A}(s)}(\theta+Z)^{\top}a\right]-\theta^{\top}\widehat{a}$. (2)
[…]
$).\mathcal{L}_{\Omega}(\theta;\widehat{a}):=\Omega^{*}(\theta)+\Omega(\widehat{a})-\langle\theta|\widehat{a}\rangle=\sup_{a\in\operatorname{dom}(\Omega)}\big(\langle\theta|a\rangle-\Omega(a)\big)-\big(\langle\theta|\widehat{a}\rangle-\Omega(\widehat{a}$$\big)$. (3)
For a given $\theta\in\operatorname{dom}(\Omega^{*})$, we introduce the regularized prediction as $\sup_{a\in\operatorname{dom}(\Omega)}\langle\theta|a\rangle-\Omega(a)$. The Fenchel-Young loss measures the non-optimality of $a\in\operatorname{dom}(\Omega)$ as a solution of the regularized prediction problem. It is nonnegative and convex in $\theta$. If in addition $\Omega$ is proper, convex, and lower semi-continuous, $\mathcal{L}_{\Omega}$ reaches zero if and only if $a$ is a solution of the regularized prediction problem.
@@
with temperature parameter $\tau$, controlling the sharpness of the softmax. If we decrease $\tau$, $\widehat{a}$ approaches the greedy action $\operatorname*{argmax}_{a^{\prime}}\left(Q_{\psi_{\beta}}(s,a^{\prime})\right)$. If we increase $\tau$, $\widehat{a}$ approaches the uniform average $\frac{1}{m}\sum_{a^{\prime}}a^{\prime}$. Note that $\widehat{a}\in\mathcal{A}(s)$ does not have to hold, as it is not executed in the environment, but only serves as a training signal for the Fenchel-Young loss.
[…]
@@
To ensure sufficient exploration of the state space $\mathcal{S}$, we introduce stochasticity into the forward pass during training: we perturb the score vector $\theta$ using Gaussian noise $Z\sim\mathcal{N}(\theta,\sigma_{f})$ with exploration standard deviation $\sigma_{f}$, sample one perturbed vector $\eta$, and select an action $a=f(\eta,s)$. Overall, we use three Gaussian perturbations of $\theta$, distinguished by their standard deviations: $\sigma_{f}$ to facilitate exploration of $\mathcal{S}$ in the forward pass, $\sigma_{b}$ to ensure exploration of the combinatorial action space $\mathcal{A}(s)$ during target action selection, and $\varepsilon$ to regularize the CO-layer $f$ for the Fenchel-Young loss.
@@ 2.2 Geometrical insights
Policies as solutions of regularized optimization over the distribution simplex
@@
A sampling-based primal-dual algorithm for the actor update
@@
Here, $\Omega^{*}$ is the Fenchel-conjugate of $\Omega$, $\nabla\Omega^{*}$ the gradient of $\Omega^{*}$, and $\partial\Omega$ the sub-differential of $\Omega$. The following proposition, proved in Appendix B, shows that the static version of Algorithm 1 can be seen as a variant of the primal-dual algorithm presented in Bouvier et al. (2025), enhancing it with an additional sampling step.
@@
where $\delta_{a}$ is the Dirac distribution on $a$, $\Omega_{\Delta}$ is the negentropy, and $\Omega_{\varepsilon,\Delta}$ is the conjugate of the sparse perturbation, both detailed in Appendix B. Since $\hat{q}_{m}^{(t+\frac{1}{2})}$ is sparse by design, we discuss the definition of gradients and sub-gradients at the (relative) boundary of the domains in Appendix B.
In Equations (8), for convenience in the implementation, we involve two distinct regularization functions on the distribution simplex $\Delta^{\mathcal{A}}$, detailed in Appendix B. The main difference between Equations (7) and Equations (8) is the sampling step for the primal update. Recall that involving the critic in Equation (7a) may be intractable. Indeed, the critic function may be highly nonlinear, and the resulting nonlinear combinatorial optimization problem may be intractable. Instead, Equation (8) is based on a sampling step, which only requires $m$ evaluations of the critic, which is tractable even when optimizing it with respect to $a\in\mathcal{A}(s)$ is not. In both primal-dual algorithms, the dual update (of $\theta$) is equivalent to solving a convex optimization problem, precisely minimizing a Fenchel-Young loss generated by $\Omega_{\mathcal{C}}$. It is very convenient in practice, since the weights $w$ of our actor policy can be updated via stochastic gradient descent, although it relies on a piecewise constant CO-layer $f$.
@@ 3 Numerical studies
We evaluate SRL across six environments that reflect industrial applications with large combinatorial action spaces. Appendix C describes the experimental setup, and Appendix D details the environments. We first consider three static environments – common industrial benchmarks from Dalle et al. (2022) – namely, a Warcraft Shortest Paths Problem, a Single Machine Scheduling Problem, and a Stochastic Vehicle Scheduling Problem. As discussed in Appendix E, SRL matches the performance of SIL, while relying solely on access to a (black-box) cost function and without requiring the expert knowledge necessary for SIL. These results underline the versatility and broader potential of SRL, even though we designed it primarily for dynamic settings. We now discuss our findings for the dynamic environments in more detail.
The dynamic environments model online decision-making in C-MDPs. We consider: i) a Dynamic Vehicle Scheduling Problem (DVSP), based on the Dynamic Vehicle Routing Problem introduced by Kool et al. (2022); Baty et al. (2024). This problem with exogenous uncertainty requires serving spatio-temporally distributed requests that are revealed over time. The goal is to find cost-minimizing routes while fulfilling all requests. ii) A Dynamic Assortment Problem (DAP), adapted from Dulac-Arnold et al. (2016) andChen et al. (2020). This problem with endogenous uncertainty involves selecting item assortments that are shown to customers, whose choices follow a multinomial logit model. Item features evolve based on past decisions, the goal is overall revenue maximization. iii) A Gridworld Shortest Paths Problem (GSPP), inspired by gridworld and robotic control tasks (Chandak et al., 2019; Zhang et al., 2020). This problem with endogenous uncertainty requires the agent to find cost-minimizing paths to targets, which move to new locations when being reached. The costs of paths are influenced by prior paths.
@@
We compare SRL against two baselines: SIL, a structured imitation learning approach, and Proximal Policy Optimization (PPO), an unstructured RL algorithm. All algorithms use identical COAML-pipelines; SRL and PPO also share the same critic architectures to ensure a fair comparison. We select SIL due to its methodological proximity to SRL and its strong performance in combinatorial settings (Baty et al., 2024; Jungel et al., 2024). We include PPO for its stability and compatibility with our pipeline architecture (Schulman et al., 2017). Alternatives such as Soft Actor-Critic would require substantial modifications to the actor and critic architectures, as would neural CO-methods (e.g., Bello et al., 2017) and Q-value-based optimization (e.g., Xu et al., 2025). To keep this analysis concise, we concentrate on comparing COAML-pipelines using different learning paradigms and leave an in-depth comparison of architectures to future work. As performance references, we include two additional baselines: an expert policy and a greedy policy. In the static DVSP, the expert has access to the complete problem instance and thus represents the offline optimum. In contrast, in the dynamic DAP and GSPP, traceability constraints limit information access, and the expert corresponds to the best possible online policy, given the sequential nature of the decision process. We train all algorithms using the same number of episodes, employing environment-specific train/validation/test splits. We tune hyperparameters per algorithm and environment, using the PPO-optimized episode numbers consistently across methods. Each algorithm is retrained using ten random seeds. Appendix C provides further details on the experimental setup and baselines.
@@
We present results for the dynamic environments in Figure 4 and Figure 5. We display the performance of final models on the train and test-datasets to measure algorithmic performance, and the development of validation rewards during training to highlight convergence behavior. While SRL performs comparably to SIL in the DVSP, it outperforms SIL in the DAP by 8% and in the GSPP by 78%, even surpassing the online optimum by 79% in the latter. In the DAP, the online optimum remains above the performance of SRL and SIL. These results highlight two key limitations of imitation learning: i) its performance is bounded by that of the expert policy; and ii) it lacks exploration to learn policies that enable escape from suboptimal states. In contrast, PPO consistently underperforms across all environments, struggling to reach greedy policies – SRL outperforms it by 16% in the DVSP, 77% in the DAP, and 92% in the GSPP. This poor performance highlights the challenges faced by unstructured RL in combinatorial action spaces. Overall, these performance gains show the superiority of SRL over SIL and PPO.
@@
[…]